# renew-dots,renew-matrix,nullify-dots

### 10 Dec 2023 v1

@@ Introduction
Complex Networked Dynamical Systems (NDS) are causally structured: the state of each unit or node evolves over time due to peer-to-peer interactions or local information exchange (Barrat, Barthélemy, and Vespignani 2008). Examples include: brain activity (Huang and Ding 2016; Liégeois et al. 2020); Gene Regulatory Networks (Aalto et al. 2020; Casadiego et al. 2017); or pandemics (Ganesh, Massoulié, and Towsley 2005; Wan et al. 2014). Understanding the underlying connectivity pattern in these applications is fundamental to forecast the long term behavior of the NDS as, for example, predicting the onset of criticality (Eroglu et al. 2020) or neurological crises (Douw et al. 2010); ascertaining whether a strain of virus persists or dies out in the long run (Lahmanovich and James 1976); or for the proper design of control or mitigation measures (Ren et al. 2019) in a pandemics.
In all these applications, usually, only the time series data of some of the nodes are readily available while the underlying causal structure linking the nodes lurks underneath. The main goal of causal inference or structure identification is to map out the underlying causal architecture of a complex system, i.e., to enable consistent inference of the latent network structure from the observed data.
@@
Reference (Machado et al. 2023) departed from this approach and addressed the framework of partial observability (with diagonal or uncorrelated noise), under a novel feature based paradigm. Instead of a scalar-based estimation for the pairs’ affinity, a feature vector is assigned to each pair of nodes, and structure is recovered by leveraging certain identifiability properties of the engineered features in a higher-dimensional space (instead of thresholding). Desirable characteristics of this feature-based approach include: i) Separability —there exists a manifold that separates the features of connected pairs from those of disconnected pairs; ii) Stability—this separation manifold should not be too sensitive to differences in the regimes of connectivity, observability and noise correlation of the underlying complex system; and iii) Locality—the feature of each pair can be computed from the time series of each pair (or neighboring pairs thereof). Separability is a necessary property to consistently cluster the features, i.e., for structure inference; stability renders supervised ML tools (e.g., SVM) amenable to generalization; and locality of the features is crucial for large-scale systems where only a subset of nodes can be feasibly observable.
However, in most applications, the underlying excitation noise exhibits nontrivial correlation structure. Depending on the regime of observation or noise structure, information about the main object of inference may be fundamentally lost (Barbier et al. 2023). The level of noise plays a major role in the feasibility of inference in that the problem can be impossible, hard (possible, but no known algorithm), or easy depending on the noise level (Barbier et al. 2019). It is thus, critical to understand the impact of the noise structure on inference problems.
@@
Our contribution is twofold: $i)$ We establish a novel condition for structural consistency of linear networked dynamical systems (Theorem $2$); $ii)$ we propose novel features to devise a Machine Learning based causal inference mechanism exhibiting competitive performance under the referred regime. Point $i)$ provides, in particular, novel guarantees for linear networked dynamical systems, where the information of the underlying network structure is contained in the observed time series (i.e., it is not fundamentally lost) and there is an algorithm to recover it. For point $ii)$, we note that the proposed feature-based causal inference method in (Machado et al. 2023) exhibits poor performance in the colored noise regime as a result of loosing the stability property and having the separability compromised (as we will formally see). We propose novel features and deploy baseline supervised methods to cluster them. The method outperforms popular estimators in the considered regime.
@@
For a recent comprehensive discussion on related causal inference works, please refer to (Machado et al. 2023) or (Matta, Santos, and Sayed 2020). Most works focus on sparse networks – by methods primarily promoting model-parsimony – and/or unstructured excitation noise, i.e., independence of the exogenous noise across distinct nodes in the system. While this renders the inference task technically tractable, for the most part, it is an overall unrealistic assumption. For example, this noise independence across nodes is a fundamental assumption in Structural Causal Models (SCM) (Pearl 2009; Peters, Janzing, and Schlkopf 2017) and underlies the structural consistency of conditional independence tests based algorithms (Spirtes 2001; Spirtes, Glymour, and Scheines 2000; Anandkumar et al. 2013; Anandkumar and Valluvan 2013) as well as other ML-based approaches (Melnychuk, Frauen, and Feuerriegel 2022; Cao et al. 2023; Runge et al. 2019; Cheng et al. 2023; Marcinkevičs and Vogt 2021). Several works explore linear networked dynamical systems (Materassi and Salapaka 2012, 2015; Pereira, Ibrahimi, and Montanari 2010; Santos, Matta, and Sayed 2020; Matta, Santos, and Sayed 2022) tailored for sparse networks and uncorrelated noise. The recent reference (Shaikh Veedu and Salapaka 2023) considers linear dynamic influence models (LDIMs), defined in the frequency domain, with correlated noise and develops a Wiener based approach tailored for sparse networks – leveraging on a sparse $+$ low rank method. The reference makes important limiting assumptions on the noise correlation: i) it results from affine interactions; and, ii) the excitation noise is assumed independent across latent nodes.
@@
Figure 1: Sufficient condition for structural consistency. The plots represent the entries of the corresponding matrices: the estimator, the ground-truth and the error.
@@
The time series data streams from the linear networked dynamical system given by
@@
where $\mathbf{y}(n)=\left[y_{1}(n)\,\,y_{2}(n)\,\,\ldots\,\,y_{N}(n)\right]^{\top}\in\mathbb{R}^{N}$ represents the state-vector of the $N$-dimensional NDS collecting the state $y_{i}(n)$ of each node $i=1,2,\ldots,N$ at time $n$; $\mathbf{x}(n)$ is zero mean input noise at time $n$ with finite first and second order moments, i.e., it admits a well-defined covariance matrix $\Sigma_{x}\in\mathbb{S}_{+}^{N}$, and $\left(\mathbf{x}(n)\right)_{n\in\mathbb{N}}$ is independent across time $n$; $A\in\mathbb{R}_{+}^{N}$ is the non-negative symmetric interaction matrix describing the graph linking the nodes as its support. We assume that the spectral radius of $A$ is smaller than $1$, $\rho(A)<1$, i.e., the dynamical system is assumed stable. Namely, we assume that $A$ is stochastic up to a multiplicative constant (smaller than $1$), i.e., $A=\rho\overline{A}$ where $\overline{A}$ is stochastic and $\rho<1$.
@@
for any given $n\in\mathbb{N}$. The $ij^{\text{th}}$ entry of the matrix $\mathcal{F}^{(n)}_{S}$ reflects the strength of the edge from $i$ to $j$ from $n$ time series samples. That is, $\mathcal{F}^{(n)}_{S}$ is a matrix summarizing the estimated affinities between the pairs of nodes in the system. The estimator $F^{(n)}_{S}$ is structurally consistent whenever the estimated strength $\mathcal{F}_{ij}^{(n)}$ of any connected pair $(i,j)$ lies above the estimated strength of any disconnected pair. In this case, the underlying network structure is recovered via proper thresholding. This is illustrated in the left-bottom of Fig. 1 and it is formalized with the following definition presented in (Machado et al. 2023).
@@
We write a matrix-valued estimator, see equation (2), as $\mathcal{F}_{S}^{(n)}=A_{S}+\mathcal{E}^{(n)}_{S}$ where, under partial observability, $A_{S}$ is the underlying ground-truth interaction matrix, $\mathcal{E}^{(n)}_{S}$ is the (matrix) error term, and $S$ is the set of observable nodes.
Example: Granger estimator. Reference (Matta, Santos, and Sayed 2020) studied the error $\mathcal{E}$ for the Granger estimator $R_{1}\left(R_{0}\right)^{-1}$, where $R_{1}$ and $R_{0}$ are the limit one-lag and correlation matrices, respectively. Under the regime of full observability, Granger recovers the underlying interaction matrix $A$ w.h.p. since $R_{1}\left(R_{0}\right)^{-1}=A$. In other words, the limiting error matrix $\mathcal{E}$ associated with Granger under full observability is zero (trivially flat). However, under partial observability, the limiting error matrix associated with Granger is given by (Matta, Santos, and Sayed 2020) $\mathcal{E}_{S}=A_{SS^{\prime}}\left(\left[R_{0}^{-1}\right]_{S^{\prime}}\right)^{-1}\left[A^{2}\right]_{S^{\prime}S}$, where $S^{\prime}$ is the set of latent nodes (complement of $S$). For diagonal and correlated noise structures, this leads to
@@
The error term $\mathcal{E}_{S}^{({\sf diagonal})}$ has been thoroughly studied in (Matta, Santos, and Sayed 2020) that shows structural consistency of the Granger estimator under partial observability and diagonal noise. However, even for Granger, there is no known analysis for $\mathcal{E}_{S}^{({\sf colored})}$ even though, in fact, under colored noise and partial observability, Granger’s performance degrades. $\square$
As we now show, for structural consistency, we do not need the error term to be zero (or to converge to zero with $n$). Define ${\sf Osc}\left(\mathcal{E}^{(n)}_{S}\right)\overset{\Delta}{=}\mathcal{E}^{(n)}_{\max}-\mathcal{E}^{(n)}_{\min}$ and let $A^{+}_{\min}$ be the smallest nonzero entry of $A_{S}$, where $\mathcal{E}^{(n)}_{\max}$ and $\mathcal{E}^{(n)}_{\min}$ are the maximum and minimum off-diagonal entries of the error matrix $\mathcal{E}_{S}^{(n)}$. Then, if, with high probability (w.h.p.),
${\sf Osc}\left(\mathcal{E}^{(n)}_{S}\right)\leq\frac{A^{+}_{\min}}{2},\vspace{-5px}$ (5)
it is easily seen that $F^{(n)}_{S}$ is structurally consistent. In fact, condition (5) precludes inversion and consequent loss of structural information in definition 1 in the following sense: If $ij$ is a connected pair and $k\ell$ is a disconnected pair, then constraint (5) implies that $\mathcal{F}_{ij}^{(n)}>\mathcal{F}^{(n)}_{k\ell}$ necessarily. This means that, under (5), structure can be recovered via properly thresholding the off-diagonal entries of the matrix $\mathcal{F}^{(n)}_{S}$, otherwise, structural information is lost. Further, when the error matrix is flat, i.e., $\mathcal{E}_{S}^{(n)}\propto\mathbf{1}\mathbf{1}^{\top}$, all the entries of $F^{(n)}$ are the entries of $A_{S}$ shifted by the same quantity. Therefore, again, structural information is preserved. Fig. 1 offers an illustration of this discussion.
In summary, the behavior of the error matrix $\mathcal{E}_{S}^{(n)}$ determines whether structural information is preserved in the observed time series or fundamentally lost: an estimator $F^{(n)}_{S}$ is structurally consistent whenever the error matrix is flat enough w.h.p. However, given a particular estimator, this error term responds differently to distinct regimes of i) connectivity of the underlying networked dynamical system, ii) observability, or iii) noise structure. This discussion shows the importance of two delicate steps in structure estimation: i) the characterization of the error term; and ii) the analysis of the flatness of the error. In other words, what matters is the variability or oscillation behavior, not the (average) drift of the entries of the error. It also alludes to the difficulty in handling the noise structure.
@@
$\Sigma_{x}:=\sigma^{2}_{{\sf gap}}I+\beta\mathbf{1}\mathbf{1}^{\top}+\overline{\Sigma}\vspace{-5px}$ (6)
where we have defined $\sigma^{2}_{{\sf gap}}\overset{\Delta}{=}\sigma^{2}-\max_{i\neq j}\mathbb{E}\left[\mathbf{x}_{i}\mathbf{x}_{j}\right]>0$ and $\beta\overset{\Delta}{=}\frac{1}{N(N-1)}\mathbf{1}^{\top}{\sf Off}(\Sigma_{x})$, with ${\sf Off}(\Sigma_{x})$ representing the vector collecting the off-diagonal entries of the matrix $\Sigma_{x}$. Thus, $\beta$ stands for the average of the diagonal entries of $\Sigma_{x}$.
@@
=$\mathcal{E}\overset{\Delta}{=}\frac{1}{\sigma_{\sf gap}^{2}}\left[\beta\rho\mathbf{1}\mathbf{1}^{\top}+\left(I-A^{2}\right)\left(\sum_{i=0}^{\infty}A^{i+1}\overline{\Sigma}A^{i}\right)\right],\vspace{-5px}$ (8)
@@
$=\beta\left(I-A^{2}\right)A\mathbf{1}\mathbf{1}^{\top}(\sum_{i=0}^{\infty}\rho^{2i})=\frac{\beta\rho(1-\rho^{2})}{(1-\rho^{2})}\mathbf{1}\mathbf{1}^{\top}\!\!=\beta\rho\mathbf{1}\mathbf{1}^{\top}$
@@
Define $R_{k}=\lim\limits_{n\rightarrow\infty}R_{k}(n)=\lim\limits_{n\rightarrow\infty}\mathbb{E}\left[\mathbf{y}(n+k)\mathbf{y}(n)^{\top}\right]$. By modifying the proof for diagonal noise $\Sigma_{x}=\sigma^{2}I$ in (Chen, Wang, and Shen 2022), we can show that, for general (spatially) correlated noise $\Sigma_{x}$, but independent across time, the empirical covariance $\widehat{R}_{1}(n)$ or $\widehat{R}_{3}(n)$ converge in probability to the exact limit covariances $R_{1}$ and $R_{3}$. ∎
Theorem 1 characterizes the error term in terms of the covariance matrix $\Sigma_{x}$. As a Corollary to Theorem $1$, if the covariance matrix $\Sigma_{x}$ is flat, i.e., $\Sigma_{x}=\sigma^{2}_{\sf gap}I+\beta\mathbf{1}\mathbf{1}^{\top}$, then the error term associated with the estimator $\widehat{R}_{1}(n)-\widehat{R}_{3}(n)$ is flat as well, which implies that the estimator is structurally consistent w.h.p. In the next section, we will characterize the set of covariance matrices $\Sigma_{x}$ that induce an error that is flat enough to imply structural consistency.
@@
$\mathcal{F}_{ij}(n)=\left(\left[R_{D}(n)\right]_{ij},\left[R_{D+1}(n)\right]_{ij},\ldots,\left[R_{M}(n)\right]_{ij}\right),\vspace{-5px}$ (11)
with $D\leq 1$ and $M\geq 3$ are affinely separable, i.e., there exists a particular hyperplane (up to a shift) that consistently partitions this set. The previous discussion translates into the feature domain (in view of Lemma $1$ in (Machado et al. 2023)): $\beta$ controls the drift of the features separating hyperplane (compromising stability), while the variation $\overline{\Sigma}$ of the off-diagonal entries of $\Sigma_{x}$ induces a spread of the features (possibly compromising separability).
To mitigate the shift in these features and to deploy supervised methods, we shall first center the features – instead of dividing by the norm of the farthest feature as done in (Machado et al. 2023). By coupling a normalized version of these features with novel ones, we will be able to further mitigate the drift and oscillation at the same time, leading to an ML method that displays competitive causal inference performance under the considered regime.
@@
In this section, we devise a novel condition on the noise structure of the NDS (1) whereby the information about the underlying network structure is contained in the observed data and can be consistently recovered (Theorem $2$). By observed data, we mean the time series collected at the subset of observed nodes $S:=\left\{m_{1},m_{2},\ldots,m_{\left|S\right|}\right\}$ in the system, namely, the sub-vector $\left[\mathbf{y}(n)\right]_{\mathcal{S}}$ over time $n\in\mathbb{N}$.
The next theorem provides a sufficient condition on the covariance matrix $\Sigma_{x}$ of the noise and the interaction matrix $A$ to lead to the linear separability and stability of the (centered) features.
@@
Property 1. [Contraction] Given a symmetric stochastic matrix $\overline{A}$, we have ${\sf Osc}\left(\overline{A}v\right)\leq{\sf Osc}\left(v\right)$, as each entry of $\overline{v}\overset{\Delta}{=}\overline{A}v$ lies in the convex hull (Hiriart-Urruty and Lemaréchal 2001) of the set $\left\{v_{1},v_{2},\ldots,v_{N}\right\}$ of the entries of the vector $v\in\mathbb{R}^{N}$ and, in particular, $\overline{v}_{i}\in\left[v_{\min},v_{\max}\right]$ for all $i$.
@@
$\frac{{\sf Osc}\left({\sf Off}(\Sigma_{x})\right)}{\sigma_{{\sf gap}}^{2}}\leq\frac{A^{+}_{\min}(1-\rho^{2})}{2\rho(\rho^{2}+1)},\vspace{-10px}$ (12)
@@
$\begin{array}[]{ccl}{\sf Osc}\left(\left(I-A^{2}\right)V\right)&\leq&{\sf Osc}\left(V\right)+{\sf Osc}\left(A^{2}V\right)\\ &\leq&{\sf Osc}(V)+\rho^{2}{\sf Osc}\left(\overline{A}^{2}V\right)\\ &\leq&\left(1+\rho^{2}\right){\sf Osc}\left(V\right)\end{array},\vspace{-5px}$ (13)
@@
$\begin{array}[]{ccl}{\sf Osc}\left(\left(I-A^{2}\right)AV\right)&=&\rho\,{\sf Osc}\left(\left(I-A^{2}\right)\overline{A}V\right)\\ &\leq&\rho\left(1+\rho^{2}\right){\sf Osc}(V).\end{array}\vspace{-5px}$ (14)
@@
$\overline{\Sigma}_{\max}\sum_{i=0}^{\infty}A^{i}\mathbf{1}\mathbf{1}^{\top}A^{i}=\frac{\overline{\Sigma}_{\max}}{1-\rho^{2}}\mathbf{1}\mathbf{1}^{\top},\vspace{-15px}$
@@
Observe that under exogenous interventions to the NDS, i.e., $\mathbf{y}(n+1)=A\mathbf{y}(n)+\mathbf{x}(n+1)+\mathbf{\xi}(n+1),$ where $\xi(n)\sim\mathcal{N}(0,\sigma_{\xi}^{2})$ is i.i.d., and independent of the process $\left(\mathbf{x}(n)\right)$, equation (12) becomes $\frac{{\sf Osc}\left({\sf Off}(\Sigma_{x})\right)}{\sigma_{{\sf gap}}^{2}+\sigma_{\xi}^{2}}\leq\frac{A_{\min}(1-\rho^{2})}{2\rho(\rho^{2}+1)},$ which implies, in particular, that regardless of the covariance matrix of the input process $\left(\mathbf{x}(n)\right)$, the features are linearly separable w.h.p. as soon as the level (or variance) $\sigma_{\xi}^{2}$ of the external intervention $\left(\xi(n)\right)$ is high enough.

### 12 Feb 2024 v3

# Learning the Causal Structure of Networked Dynamical Systems under Latent Nodes and Structured Noise

@@ Introduction
Complex Networked Dynamical Systems (NDS) are causally structured: the state of each unit or node evolves over time due to peer-to-peer interactions or local information exchange (Barrat, Barthélemy, and Vespignani 2008). Examples include: brain activity (Huang and Ding 2016; Liégeois et al. 2020); Gene Regulatory Networks (Aalto et al. 2020; Casadiego et al. 2017); or pandemics (Ganesh, Massoulié, and Towsley 2005; Wan et al. 2014). Understanding the underlying connectivity pattern in these applications is fundamental to forecast the long term behavior of the NDS as, for example, predicting the onset of criticality (Eroglu et al. 2020) or neurological crises (Douw et al. 2010); ascertaining whether a strain of virus persists or dies out in the long run (Lahmanovich and James 1976); or for structure-informed downstream tasks, e.g., the design of mitigation measures (Ren et al. 2019) in a pandemics.
In all these applications, usually, only the time series data at some of the nodes are readily available while the underlying causal structure linking these nodes lurks underneath. The main goal of causal inference or structure identification is to map out the underlying causal architecture of a complex system, i.e., to enable consistent inference of the latent network structure from the observed data.
@@
Reference (Machado et al. 2023) departed from this approach and addressed the framework of partial observability (with diagonal or uncorrelated noise), under a novel feature based paradigm. Instead of a scalar-based estimation for the pairs’ affinity, a feature vector is assigned to each pair of nodes, and structure is recovered by leveraging certain identifiability properties of the engineered features in a higher-dimensional space (instead of thresholding). Desirable characteristics of this feature based approach include: $i)$ Separability —there exists a manifold that separates the features of connected pairs from those of disconnected pairs; $ii)$ Stability—this separation manifold should not be too sensitive to differences in the regimes of connectivity, observability and noise correlation of the underlying complex system; and $iii)$ Locality—the feature of each pair can be computed from the time series of each pair (or neighboring pairs thereof). Separability is a necessary property to consistently cluster the features, i.e., for structure inference; stability renders supervised ML tools (e.g., SVM) amenable to generalization; and locality of the features is crucial for large-scale systems where only a subset of nodes can be feasibly observable.
However, in most applications, the underlying excitation noise exhibits nontrivial correlation structure. Depending on the regime of observation or noise structure, information about the main object of inference may be fundamentally lost (Barbier et al. 2023). The level of noise plays a major role in the feasibility of statistical inference tasks in that the problem can be impossible, hard (possible, but no known polynomial-time algorithm), or easy, depending on the noise level (Barbier et al. 2019). It is thus, critical to understand the impact of the noise structure on inference problems.
@@
Our contribution is twofold: $i)$ we establish a novel identifiability condition for linear networked dynamical systems (Theorem $2$); $ii)$ we propose novel features to devise a Machine Learning based causal inference mechanism exhibiting competitive performance under the referred regime. Point $i)$ provides, in particular, novel guarantees for linear networked dynamical systems, where the information of the underlying network structure is contained in the observed time series data (i.e., it is not fundamentally lost).
@@
For a recent comprehensive discussion on related causal inference works, please refer to (Machado et al. 2023) or (Matta, Santos, and Sayed 2020). Most works focus on sparse networks – by methods primarily promoting model-parsimony – and/or unstructured excitation noise, i.e., independence of the exogenous noise across distinct nodes in the system. While this renders the inference task technically tractable, for the most part, it is an overall unrealistic assumption. For example, this noise independence across nodes is a fundamental assumption in Structural Causal Models (SCM) (Pearl 2009; Peters, Janzing, and Schlkopf 2017) and underlies the structural consistency of conditional independence tests based algorithms (Spirtes 2001; Spirtes, Glymour, and Scheines 2000; Anandkumar et al. 2013; Anandkumar and Valluvan 2013) as well as other ML-based approaches (Melnychuk, Frauen, and Feuerriegel 2022; Cao et al. 2023; Runge et al. 2019; Cheng et al. 2023; Marcinkevičs and Vogt 2021). Several works explore linear networked dynamical systems (Materassi and Salapaka 2012, 2015; Pereira, Ibrahimi, and Montanari 2010; Santos, Matta, and Sayed 2020; Matta, Santos, and Sayed 2022) tailored to sparse networks and uncorrelated noise. The recent reference (Shaikh Veedu and Salapaka 2023) considers linear dynamic influence models (LDIMs), defined in the frequency domain, with correlated noise and develops a Wiener based approach specialized to sparse networks – leveraging on a sparse $+$ low rank method. The reference makes important limiting assumptions on the noise correlation: $i)$ it results from affine interactions; and, $ii)$ the excitation noise is assumed independent across latent nodes.
@@
Figure 1: Sufficient condition for structural consistency. The plots represent the off-diagonal entries of the corresponding matrices: the estimator, the ground truth and the error.
@@
The time series data streams from the linear Networked Dynamical System (NDS) given by
@@
where $\mathbf{y}(n)=\left[y_{1}(n)\,\,y_{2}(n)\,\,\ldots\,\,y_{N}(n)\right]^{\top}\in\mathbb{R}^{N}$ represents the state-vector of the $N$-dimensional NDS collecting the state $y_{i}(n)$ of each node $i=1,2,\ldots,N$ at time $n$; $\mathbf{x}(n)$ is the zero mean input noise at time $n$ with finite first and second order moments, i.e., it admits a well-defined covariance matrix $\Sigma_{x}\in\mathbb{S}_{+}^{N}$, and $\left(\mathbf{x}(n)\right)_{n\in\mathbb{N}}$ is independent across time $n$; $A\in\mathbb{R}_{+}^{N}$ is the non-negative symmetric interaction matrix describing the graph linking the nodes as its support. We assume that the spectral radius of $A$ is smaller than $1$, $\rho(A)<1$, i.e., the dynamical system is assumed stable. Namely, we assume that $A$ is stochastic up to a multiplicative constant, i.e., $A=\rho\overline{A}$ where $\overline{A}$ is stochastic and $\rho<1$.
@@
for any given $n\in\mathbb{N}$. The $ij^{\text{th}}$ entry of the matrix $\mathcal{F}^{(n)}_{S}$ reflects the strength of the edge from $j$ to $i$ from $n$ time series samples. That is, $\mathcal{F}^{(n)}_{S}$ is a matrix summarizing the estimated affinities between the pairs of nodes in the system. The estimator $F^{(n)}_{S}$ is structurally consistent whenever the estimated strength $\mathcal{F}_{ij}^{(n)}$ of any connected pair $(i,j)$ lies above the estimated strength of any disconnected pair. In this case, the underlying network structure is recovered via proper thresholding. This is illustrated in Fig. 1 and it is formalized with the following definition presented in (Machado et al. 2023).
@@
We write a matrix-valued estimator, see equation (2), as $\mathcal{F}_{S}^{(n)}=A_{S}+\mathcal{E}^{(n)}_{S}$ where, under partial observability, $A_{S}$ is the underlying ground truth interaction matrix, $\mathcal{E}^{(n)}_{S}$ is the (matrix) error term, and $S$ is the set of observable nodes. As we now show, for structural consistency, we do not need the error term to be zero (or to converge to zero with $n$). Define ${\sf Osc}\left(\mathcal{E}^{(n)}_{S}\right)\overset{\Delta}{=}\mathcal{E}^{(n)}_{\max}-\mathcal{E}^{(n)}_{\min}$ and let $A^{+}_{\min}$ be the smallest nonzero entry of $A_{S}$, where $\mathcal{E}^{(n)}_{\max}$ and $\mathcal{E}^{(n)}_{\min}$ are the maximum and minimum off-diagonal entries of the error matrix $\mathcal{E}_{S}^{(n)}$. Then, if, with high probability (w.h.p.),
${\sf Osc}\left(\mathcal{E}^{(n)}_{S}\right)\leq\frac{A^{+}_{\min}}{2},$ (4)
it is easily seen that $F^{(n)}_{S}$ is structurally consistent. In fact, condition (4) precludes inversion and consequent loss of structural information in Definition 1 in the following sense: If $ij$ is a connected pair and $k\ell$ is a disconnected pair, then constraint (4) implies that $\mathcal{F}_{ij}^{(n)}>\mathcal{F}^{(n)}_{k\ell}$ necessarily. This means that, under (4), structure can be recovered via properly thresholding the off-diagonal entries of the matrix $\mathcal{F}^{(n)}_{S}$, otherwise, structural information is lost. Further, when the error matrix is flat, i.e., ${\sf Off}\left(\mathcal{E}_{S}^{(n)}\right)\propto{\sf Off}\left(\mathbf{1}\mathbf{1}^{\top}\right)$, all the entries of $F^{(n)}$ are the entries of $A_{S}$ shifted by the same quantity. Therefore, again, structural information is preserved. Fig. 1 offers an illustration of this discussion.
Example: Granger estimator. Reference (Matta, Santos, and Sayed 2020) studied the error $\mathcal{E}_{S}$ for the Granger estimator $\left[R_{1}\right]_{S}\left(\left[R_{0}\right]_{S}\right)^{-1}$, under partial observability, where $R_{1}$ and $R_{0}$ are the limit one-lag and correlation matrices, respectively. Under full observability, Granger recovers the underlying interaction matrix $A$ w.h.p. since $R_{1}\left(R_{0}\right)^{-1}=A$. In other words, the limiting error matrix $\mathcal{E}$ associated with Granger under full observability is zero (trivially flat). However, under partial observability, the limiting error matrix associated with Granger is given by (Matta, Santos, and Sayed 2020) $\mathcal{E}_{S}=A_{SS^{\prime}}\left(\left[R_{0}^{-1}\right]_{S^{\prime}}\right)^{-1}\left[A^{2}\right]_{S^{\prime}S}$, where $S^{\prime}$ is the set of latent nodes (complement of $S$). For diagonal and correlated noise structures, this leads to
@@
The error $\mathcal{E}_{S}^{({\sf diagonal})}$ has been thoroughly studied in (Matta, Santos, and Sayed 2020) that shows structural consistency of the Granger estimator under partial observability and diagonal noise. However, even for Granger, there is no known analysis of $\mathcal{E}_{S}^{({\sf colored})}$. In fact, under colored noise and partial observability, Granger’s performance degrades. $\square$
In summary, the behavior of the error matrix $\mathcal{E}_{S}^{(n)}$ determines whether structural information is preserved in the observed time series or fundamentally lost: an estimator $F^{(n)}_{S}$ is structurally consistent whenever the error matrix is flat enough w.h.p. However, given a particular estimator, this error term responds differently to distinct regimes of $i)$ connectivity of the underlying networked dynamical system, $ii)$ observability, or $iii)$ noise structure. This discussion shows the importance of two delicate steps in structure estimation: $i)$ the characterization of the error term; and $ii)$ the analysis of the flatness of the error. Indeed, what matters is the variability and not the (average) drift of its entries. It also alludes to the difficulty in handling the noise structure.
@@
$\Sigma_{x}:=\sigma^{2}_{{\sf gap}}I+\beta\mathbf{1}\mathbf{1}^{\top}+\overline{\Sigma}$ (6)
where we have defined $\sigma^{2}_{{\sf gap}}\overset{\Delta}{=}\sigma^{2}-\max_{i\neq j}\mathbb{E}\left[\mathbf{x}_{i}\mathbf{x}_{j}\right]>0$ and $\beta\overset{\Delta}{=}\frac{1}{N(N-1)}\mathbf{1}^{\top}{\sf Off}(\Sigma_{x})$, with ${\sf Off}(\Sigma_{x})$ representing the vector collecting the off-diagonal entries of the matrix $\Sigma_{x}$. Thus, $\beta$ stands for the average of the off-diagonals of $\Sigma_{x}$.
@@
=$\mathcal{E}\overset{\Delta}{=}\frac{1}{\sigma_{\sf gap}^{2}}\left[\beta\rho\mathbf{1}\mathbf{1}^{\top}+\left(I-A^{2}\right)\left(\sum_{i=0}^{\infty}A^{i+1}\overline{\Sigma}A^{i}\right)\right],$ (8)
@@
$=\beta\left(I-A^{2}\right)A\mathbf{1}\mathbf{1}^{\top}(\sum_{i=0}^{\infty}\rho^{2i})=\frac{\beta\rho(1-\rho^{2})}{(1-\rho^{2})}\mathbf{1}\mathbf{1}^{\top}\!\!=\beta\rho\mathbf{1}\mathbf{1}^{\top}.$
@@
Define $R_{k}=\lim\limits_{n\rightarrow\infty}R_{k}(n)=\lim\limits_{n\rightarrow\infty}\mathbb{E}\left[\mathbf{y}(n+k)\mathbf{y}(n)^{\top}\right]$. By modifying the proof for diagonal noise $\Sigma_{x}=\sigma^{2}I$ in (Chen, Wang, and Shen 2022), we can show that, for general (spatially) correlated noise $\Sigma_{x}$, but independent across time, the empirical covariances $\widehat{R}_{1}(n)$ or $\widehat{R}_{3}(n)$ converge in probability to the exact limit covariances $R_{1}$ and $R_{3}$. ∎
Theorem 1 characterizes the error term in terms of the covariance matrix $\Sigma_{x}$. As a Corollary to Theorem $1$, if $\Sigma_{x}$ is flat, i.e., $\Sigma_{x}=\sigma^{2}_{\sf gap}I+\beta\mathbf{1}\mathbf{1}^{\top}$, then the error term associated with the estimator $\widehat{R}_{1}(n)-\widehat{R}_{3}(n)$ is flat as well, implying that the estimator is structurally consistent w.h.p.
@@
$\mathcal{F}_{ij}(n)=\left(\left[R_{D}(n)\right]_{ij},\left[R_{D+1}(n)\right]_{ij},\ldots,\left[R_{M}(n)\right]_{ij}\right),$ (11)
with $D\leq 1$ and $M\geq 3$ are affinely separable, i.e., there exists a particular hyperplane (up to a shift) that consistently partitions this set. The previous discussion translates into the feature domain (in view of Lemma $1$ in (Machado et al. 2023)): $\beta$ controls the drift of the features separating hyperplane (compromising stability), while the variation $\overline{\Sigma}$ of the off-diagonal entries of $\Sigma_{x}$ induces a spread of the features.
@@
The next theorem provides a sufficient condition on the noise covariance $\Sigma_{x}$ and the interaction matrix $A$ that lead to the linear separability and stability of the (centered) features.
@@
Property 1. [Contraction] Given a symmetric stochastic matrix $\overline{A}$, we have ${\sf Osc}\left(\overline{A}v\right)\leq{\sf Osc}\left(v\right)$, as each entry of $\overline{v}\overset{\Delta}{=}\overline{A}v$ lies in the convex hull (Hiriart-Urruty and Lemaréchal 2001) of the set $\left\{v_{1},v_{2},\ldots,v_{N}\right\}$ of the entries of $v\in\mathbb{R}^{N}$ and, in particular, $\overline{v}_{i}\in\left[v_{\min},v_{\max}\right]$ for all $i$.
@@
$\frac{{\sf Osc}\left({\sf Off}(\Sigma_{x})\right)}{\sigma_{{\sf gap}}^{2}}\leq\frac{A^{+}_{\min}(1-\rho^{2})}{2\rho(\rho^{2}+1)},$ (12)
@@
$\begin{array}[]{ccl}{\sf Osc}\left(\left(I-A^{2}\right)V\right)&\leq&{\sf Osc}\left(V\right)+{\sf Osc}\left(A^{2}V\right)\\ &\leq&{\sf Osc}(V)+\rho^{2}{\sf Osc}\left(\overline{A}^{2}V\right)\\ &\leq&\left(1+\rho^{2}\right){\sf Osc}\left(V\right)\end{array},$ (13)
@@
$\begin{array}[]{ccl}{\sf Osc}\left(\left(I-A^{2}\right)AV\right)&=&\rho\,{\sf Osc}\left(\left(I-A^{2}\right)\overline{A}V\right)\\ &\leq&\rho\left(1+\rho^{2}\right){\sf Osc}(V).\end{array}$ (14)
@@
$\overline{\Sigma}_{\max}\sum_{i=0}^{\infty}A^{i}\mathbf{1}\mathbf{1}^{\top}A^{i}=\frac{\overline{\Sigma}_{\max}}{1-\rho^{2}}\mathbf{1}\mathbf{1}^{\top},$
@@
Under exogenous interventions to the NDS, i.e., $\mathbf{y}(n+1)=A\mathbf{y}(n)+\mathbf{x}(n+1)+\mathbf{\xi}(n+1),$ where $\xi(n)\sim\mathcal{N}(0,\sigma_{\xi}^{2})$ is i.i.d., and independent of the process $\left(\mathbf{x}(n)\right)$, equation (12) becomes $\frac{{\sf Osc}\left({\sf Off}(\Sigma_{x})\right)}{\sigma_{{\sf gap}}^{2}+\sigma_{\xi}^{2}}\leq\frac{A_{\min}(1-\rho^{2})}{2\rho(\rho^{2}+1)},$ implying that regardless of the covariance matrix of the input process $\left(\mathbf{x}(n)\right)$, the features are linearly separable w.h.p. as soon as the level (or variance) $\sigma_{\xi}^{2}$ of the external intervention $\left(\xi(n)\right)$ is high enough.